# The multiplicative Kowalski-Słodkowski Theorem for Hermitian algebras

### v1

@@ 3 The multiplicative Kowalski-Słodkowski Theorem for Hermitian algebras
Let $\left(a_{n}\right)$ and $\left(b_{n}\right)$ be sequences of positive elements in $A$ such that $a_{n}^{2}<b_{n}^{2}$ for all $n\in\mathbb{N}$. If $\left(v_{n}\right)$ is a sequence in $A$ such that $s(b_{n}v_{n})\to 0$, then $s(a_{n}v_{n})\to 0$.
@@
Since $a_{n}^{2}<b_{n}^{2}$ for every $n$, it follows that $v_{n}^{\star}a_{n}^{2}v_{n}\leq v_{n}^{\star}b_{n}^{2}v_{n}$ for each $n$. Moreover, $v_{n}^{\star}b_{n}^{2}v_{n}\leq\rho(v_{n}^{\star}b_{n}^{2}v_{n})\mathbf{1}$ and hence $\rho(v_{n}^{\star}a_{n}^{2}v_{n})\leq\rho(v_{n}^{\star}b_{n}^{2}v_{n}),$ for every $n\in\mathbb{N}.$ But then,
@@
Let $0\leq a$ and let $(v_{n})$ be a sequence in $A$ such that $s\left(\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}\,v_{n}\right)\to 0$. Then $s(av_{n})\to 0$.
@@
Observe first that, by hypothesis, $s\left(2\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}\,v_{n}\right)\to 0$ and that for any $n\in\mathbb{N}$
$0<\left(\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}-a\right)^{2}<\left(2\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}\right)^{2}.$
Hence, Lemma 3.1 says that $s\left(\left(\sqrt{a+\frac{1}{n^{2}}\mathbf{1}}-a\right)v_{n}\right)\to 0$. Using the subadditivity of $s$, we get that $s(av_{n})\to 0$. ∎

### v2

@@ 3 The multiplicative Kowalski-Słodkowski Theorem for Hermitian algebras
Let $\left(a_{n}\right)$ and $\left(b_{n}\right)$ be sequences in $S_{A}$ such that $a_{n}^{2}<b_{n}^{2}$ for all $n\in\mathbb{N}$. If $\left(v_{n}\right)$ is a sequence in $A$ such that $s(b_{n}v_{n})\to 0$, then $s(a_{n}v_{n})\to 0$.
@@
Since $0<b_{n}^{2}-a_{n}^{2}$ it follows, using the Shirali-Ford Theorem, that $v_{n}^{\star}a_{n}^{2}v_{n}\leq v_{n}^{\star}b_{n}^{2}v_{n}$ for each $n$. Moreover, $v_{n}^{\star}b_{n}^{2}v_{n}\leq\rho(v_{n}^{\star}b_{n}^{2}v_{n})\mathbf{1}$ and hence $\rho(v_{n}^{\star}a_{n}^{2}v_{n})\leq\rho(v_{n}^{\star}b_{n}^{2}v_{n}),$ for every $n\in\mathbb{N}.$ But then,
@@
Let $a\in S_{A}$, and let $(v_{n})$ be a sequence in $A$ such that $s\left(\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}\,v_{n}\right)\to 0$. Then $s(av_{n})\to 0$.
@@
Obviously, from the assumption, $s\left(2\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}\,v_{n}\right)\to 0$. Observe that for each $n\in\mathbb{N}$
$0<\left(\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}-a\right)^{2}<\left(2\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}\right)^{2}.$
Hence, Lemma 3.1 says that $s\left(\left(\sqrt{a^{2}+\frac{1}{n^{2}}\mathbf{1}}-a\right)v_{n}\right)\to 0$. Using the subadditivity of $s$, we get that $s(av_{n})\to 0$. ∎